# Evolving Hard Maximum Cut Instances for Quantum Approximate Optimization Algorithms

### v1

@@ 5.2. Results
5.2.1. Evaluation of the GAE
During the training phase, we randomly save 60 graphs out of 6,000 in each epoch, accumulating a total of 6,000 graphs for in-depth evaluation. Our initial examination focuses on the 64-dimensional latent space structure, utilizing embeddings generated from these 6,000 graphs. Initially, we compute several key features, as detailed in Section 7.1 (see the supplementary material (Authors, s)), prioritizing those that significantly influence the complexity and difficulty of the maximum cut problem. Research indicates that certain graph characteristics, especially eigenvalues, provide substantial insight into a graph’s connectivity and complexity concerning the maximum cut problem (Dunning et al., 2018; Moussa et al., 2020). Figure 2 presents the second-largest eigenvalues of the normalized graph Laplacian for 20-node and 100-node graphs. Nodes represent graph embedding clusters identified via DBSCAN clustering (Ester et al., 1996; Schubert et al., 2017), with edges indicating feature overlap or transitions between clusters. Node color reflects the second-largest eigenvalue, revealing graph connectivity and structural complexity. The visualization demonstrates the smooth consistency of embeddings, validating the well-trained GAE’s suitability for generating new graphs for further analysis.
We continue to evaluate our trained GAE through a link prediction task, leveraging learned embeddings to capture graph topology. These embeddings facilitate adjacency matrix reconstruction, enabling predictions of potential node connections. Performance is measured using the macro area under the receiver operating characteristics curve (ROC-AUC) (0 to 1 scale, where 0.5 indicates random guessing and 1 denotes perfect prediction). Our initial assessment focuses on a GAE trained on 10-node Erdős–Rényi graphs (50% probability) using a single A100 GPU with a batch size of 32. This setup yields a macro ROC-AUC of $0.933\pm 0.004$ across 6,000 graphs training graphs. Building on this promising result, we extend our analysis to larger graphs containing all graph models, obtaining macro ROC-AUC scores of $0.709\pm 0.001$ for 20-node graphs and $0.611\pm 0.001$ for 100-node graphs, using the same number of training samples.
5.2.2. Feature analysis through parallel coordinate plot
@@
To ensure diversity, each CMA-ES run begins from different latent points within the latent space. For 20-node maximum cut instances, we generate 1,519 instances where RQAOA outperformed GW and 1,231 instances where GW exceeded the performance of RQAOA. Similarly, for 100-node instances, both algorithms outperform each other in 850 cases. To further analyze these instances, we examine various features (see section 7.1 in the supplementary material (Authors, s)). Comparative results are visualized in Figure 3(a) for 20-node graphs and Figure 3(b) for 100-node graphs.
@@
Figure 3(a) presents feature comparisons for maximum cut instances where RQAOA and GW exhibit differing performance. The top subplot highlights cases where RQAOA outperforms GW, with darker shades indicating greater superiority. The bottom subplot shows instances favoring GW, revealing that lower performance ratios (indicating stronger performance) appear more frequently when GW surpasses RQAOA. RQAOA-dominant instances exhibit a narrow performance ratio range (0.952–0.979), whereas GW-favored cases span a wider range (0.720–0.990). A key distinction emerges in LOG_SMALLESTEIGVAL, where RQAOA-preferred graphs fall below -20, suggesting structural isolation. A representative case, featuring a mostly connected graph with one isolated node, illustrates this trend. Running 100 trials per algorithm on this instance, RQAOA outperforms GW in only 4 trials, with identical results in the rest, implying that GW’s inherent randomness may influence its comparative performance. Figure 3(b) reveals a similar trend for 100-node graphs, indicating no significant differences in computed features compared to the 20-node case.
@@
5.2.3. Feature analysis with machine learning pipeline
@@
Table 1. Important features with the classifier built on 20-node maximum cut instances.
@@
Table 2. Important features with the classifier built on 100-node maximum cut instances.
To analyze the GW algorithm’s role in generating hard maximum cut instances, we expand our feature set based on prior research (Moussa et al., 2020). Incorporating six additional GW-related features (see Section 7.2 in the supplementary material (Authors, s)), we increase our total to 18 features within our machine learning pipeline. NP-hard features like the domination number are considered but excluded due to their computational complexity and minimal impact on model accuracy. For classification, we use the TPOP library (Olson et al., 2016), an AutoML tool for optimizing machine learning pipelines. Instances, where RQAOA outperforms GW, are labeled 0, while GW-dominant cases are labeled 1. We split the dataset 70:30 for training and testing. Due to class imbalance, particularly in 20-node experiments, we optimize for balanced accuracy (Brodersen et al., 2010), while standard accuracy (Pedregosa et al., 2011) is used for the balanced 100-node dataset. Using TPOP, we conduct 10-fold cross-validation over 100 generations, evaluating 60 candidate solutions per generation.
In our 20-node experiments, we achieve an outstanding average balanced accuracy of $0.9988\pm 0.0012$ across ten runs, each with a unique random data split. This result is obtained using an ExtraTreesClassifier (Geurts et al., 2006) (parameters detailed in section 7.3). Key predictive features include EXPECTED_COSTGW_OVER_SDP_COST and SPECTRAL_GAP, where the former evaluates GW’s effectiveness, and the latter reflects graph spectral properties as shown in Table 1. The relevance of the GW algorithm’s output in our analysis aligns with our focus on instances that are particularly challenging for maximum cut problems, influenced by the GW performance. The importance of eigenvalues also plays a crucial role, as seen in LOG_SECONDLARGESTEIGVAL and LOG_LARGESTEIGVAL, while GIRTH (shortest cycle length) contributes little to classifier performance. Further, the partial dependence plot visualizes the impact of two top features as shown in Figure 4(a). This plot demonstrates how these features influence the probability of predicting class 1, which in this study corresponds to instances where GW outperforms RQAOA. The plot reveals a clear trend: as the feature value changes, the probability of an instance being classified as class 1 increases, indicating a positive correlation between these features and the probability of challenging GW.
[…]
In the analysis of 100-node instances, we achieve perfect accuracy, recorded as $1.0000\pm 0.0000$, using an ExtraTreesClassifier. The parameters for this classifier are outlined in section 7.3. As indicated in Table 2, the feature EXPECTED_COSTGW_OVER_SDP_COST stands out as the most significant, showing a much higher influence compared to other features. This finding is in line with our earlier results from the 20-node experiments, where four out of the five most influential features are similarly linked to the performance metrics of the GW algorithm. Consistent with previous observations, the feature GIRTH, which do not impact the classifier’s decisions. The influence of the primary features is visualized in Figure 4(b) through a partial dependence plot. It illustrates that as the values of EXPECTED_COSTGW_OVER_SDP_COST and LOG_SMALLESTEIGVAL increase, there is a corresponding rise in the probability that an instance will perform better under GW compared to RQAOA.
To assess the significance of features, we conduct a permutation important analysis. This involved randomly shuffling the values within columns of our dataset, focusing particularly on the top two features identified in both the 20-node and 100-node experiments. For each feature, we generate 10 different permutations of our original dataset. We then evaluate the impact of these permutations on the model’s performance, which is initially trained on the unaltered dataset, using the same 10 seeds as the original setup. For the 20-node experiments, the permutation of the feature EXPECTED_COSTGW_OVER_SDP_COST results in a slight decrease in performance, with metrics changing by a $-0.0006\pm 0.0018$. Interstingly, shuffling the SPECTRAL_GAP feature leads to a minor performance increase, indicated by $0.0001\pm 0.0017$. When both features are permuted simultaneously, the combined effect causes a more noticeable performance decrease of $0.0009\pm 0.00177$. In the 100-node experiments, permutating the EXPECTED_COSTGW_OVER_SDP_COST feature lead to a more pronounced decline in performance, measured at $-0.0067\pm 0.0048$. Permutation of the feature LOG_SMALLEST
EIGVAL results in a smaller drop in performance, recorded as $-0.0004$
±$\pm 0.0008$ on performance due to the. Permuting both features together causes a decrease in performance of $-0.0012\pm 0.0013$. Overall, we can conclude that the performance shifts upon permuting a single feature or two top features slightly affect the model’s performance.
5.2.4. RQAOA as a high-performing heuristic with machine leaning pipeline
(a) Partial dependence plot highlighting key features that distinguish significant RQAOA performance gains on 20-node maximum cut instances, based on permutation importance.
(b) Partial dependence plot highlighting key features that distinguish significant RQAOA performance gains on 100-node maximum cut instances, based on permutation importance.
[…]

### v2

@@ 5.2. Results
Evaluation of the GAE
During the training phase, we randomly save 60 graphs from a different set of 6 000 graphs in each epoch, accumulating a total of 6 000 graphs for in-depth evaluation. Our initial examination focuses on the 64-dimensional latent space structure, utilizing embeddings generated from these 6 000 graphs. Initially, we compute several key features, as detailed in section A (see the supplementary material (Pan et al., 2025)), prioritizing those that significantly influence the complexity and difficulty of the maximum cut problem. Research indicates that certain graph characteristics, especially eigenvalues, provide substantial insight into a graph’s connectivity and complexity concerning the maximum cut problem (Dunning et al., 2018; Moussa et al., 2020).
Fig. 3 presents the log-transformed second-largest eigenvalues of the normalized Laplacian for both 20-node and 100-node graph instances. Each node in the plot corresponds to a cluster of graph embeddings identified through DBSCAN (Ester et al., 1996; Schubert et al., 2017), with edges indicating transitions or similarities between clusters in the latent space. We focus on the second-largest eigenvalue, rather than others, such as the spectral gap, because it provides more nuanced insight into structural properties such as loosely connected regions or modular communities, particularly when the largest eigenvalue is close to its upper bound, making the gap less informative. The color gradient across nodes highlights continuous changes in this spectral feature, reflecting varying levels of connectivity and complexity across clusters. This smooth variation demonstrates the smooth consistency of embeddings, validating the well-trained GAE’s suitability for generating new graphs for further analysis.
We continue to evaluate our trained GAE through a link prediction task, leveraging learned embeddings to capture graph topology. These embeddings facilitate adjacency matrix reconstruction, enabling predictions of potential node connections. Performance is measured using the macro area under the receiver operating characteristics curve (ROC-AUC) (0 to 1 scale, where 0.5 indicates random guessing and 1 denotes perfect prediction). Our initial assessment focuses on a GAE trained on 10-node Erdős–Rényi graphs (50% probability) using a single A100 GPU with a batch size of 32. This setup yields a macro ROC-AUC of $0.933\pm 0.004$ across 6 000 graphs training graphs. Building on this promising result, we extend our analysis to larger graphs containing all graph models, obtaining macro ROC-AUC scores of $0.709\pm 0.001$ for 20-node graphs and $0.611\pm 0.001$ for 100-node graphs, using the same number of training samples.
Feature analysis through parallel coordinate plot
@@
To ensure diversity, each CMA-ES run begins from different latent points within the latent space. For 20-node maximum cut instances, we generate 1 519 instances where RQAOA outperformed GW and 1 231 instances where GW exceeded the performance of RQAOA. Similarly, for 100-node instances, both algorithms outperform each other in 850 cases. To further analyze these instances, we examine various features (see section A in the supplementary material (Pan et al., 2025)). Comparative results are visualized in Fig. 4(a) for 20-node graphs and Fig. 4(b) for 100-node graphs.
@@
Fig. 4(a) presents a parallel coordinate plot comparing graph features for instances where either RQAOA or GW outperforms the other. Each line in the plot corresponds to a single graph instance, with the vertical position indicating the value of each feature for that instance. The top subplot highlights cases where RQAOA outperforms GW, with darker shades indicating greater superiority. The bottom subplot shows instances favoring GW, revealing that lower performance ratios (indicating stronger performance) appear more frequently when GW surpasses RQAOA. RQAOA-dominant instances exhibit a narrow performance ratio range (0.952–0.979), whereas GW-favored cases span a wider range (0.720–0.990). A key distinction emerges in LOG_SMALLESTEIGVAL, where RQAOA-preferred graphs fall below -20, suggesting structural isolation. A representative case, featuring a mostly connected graph with one isolated node, illustrates this trend. Running 100 trials per algorithm on this instance, RQAOA outperforms GW in only 4 trials, with identical results in the rest, implying that GW’s inherent randomness may influence its comparative performance. Fig. 4(b) reveals a similar trend for 100-node graphs, indicating no significant differences in computed features compared to the 20-node case.
Feature analysis with machine learning pipeline
To analyze the GW algorithm’s role in generating hard maximum cut instances, we expand our feature set based on prior research (Moussa et al., 2020). Incorporating six additional GW-related features (see section B in the supplementary material (Pan et al., 2025)), we increase our total to 18 features within our machine learning pipeline. NP-hard features like the domination number are considered but excluded due to their computational complexity and minimal impact on model accuracy.
In our 20-node experiments, we achieve an outstanding average balanced accuracy of $0.9988\pm 0.0012$ across ten runs, each with a unique random data split. This result is obtained using an ExtraTreesClassifier (Geurts et al., 2006) (parameters detailed in section C in the supplementary material (Pan et al., 2025)). Key predictive features include EXPECTED_COSTGW_OVER_SDP_COST and SPECTRAL_GAP, where the former evaluates GW’s effectiveness, and the latter reflects graph spectral properties as shown in Table A1 (see the supplementary material (Pan et al., 2025)). The relevance of the GW algorithm’s output in our analysis aligns with our focus on instances that are particularly challenging for maximum cut problems, influenced by the GW performance. The importance of eigenvalues also plays a crucial role, as seen in LOG_SECONDLARGESTEIGVAL and LOG_LARGESTEIGVAL, while GIRTH (shortest cycle length) contributes little to classifier performance. Further, the partial dependence plot visualizes the impact of two top features as shown in Fig. 5(a). This plot demonstrates how these features influence the probability of predicting class 1, which in this study corresponds to instances where GW outperforms RQAOA. The plot reveals a clear trend: as the feature value changes, the probability of an instance being classified as class 1 increases, indicating a positive correlation between these features and the probability of challenging GW.
[…]
In the analysis of 100-node instances, we achieve perfect accuracy, recorded as $1.0000\pm 0.0000$, using an ExtraTreesClassifier. The parameters for this classifier are outlined in section C (see the supplementary material (Pan et al., 2025)). As indicated in Table A2 (see the supplementary material (Pan et al., 2025)), the feature EXPECTED_COSTGW_OVER_SDP
_COST stands out as the most significant, showing a much higher influence compared to other features. This finding is in line with our earlier results from the 20-node experiments, where four out of the five most influential features are similarly linked to the performance metrics of the GW algorithm. Consistent with previous observations, the feature GIRTH, which do not impact the classifier’s decisions. The influence of the primary features is visualized in Fig. 5(b) through a partial dependence plot. It illustrates that as the values of EXPECTED_COSTGW_OVER_SDP_COST and LOG_SMALLESTEIGVAL increase, there is a corresponding rise in the probability that an instance will perform better under GW compared to RQAOA.
To assess the significance of features, we further conduct a permutation important analysis. This involved randomly shuffling the values within columns of our dataset, focusing particularly on the top two features identified in both the 20-node and 100-node experiments. For each feature, we generate 10 different permutations of our original dataset. We then evaluate the impact of these permutations on the model’s performance, which is initially trained on the unaltered dataset, using the same 10 seeds as the original setup. For the 20-node experiments, the permutation of the feature EXPECTED_COSTGW_OVER_SDP_COST results in a slight decrease in performance, with metrics changing by a $-0.0006\pm 0.0018$. Interstingly, shuffling the SPECTRAL_GAP feature leads to a minor performance increase, indicated by $0.0001\pm 0.0017$. When both features are permuted simultaneously, the combined effect causes a more noticeable performance decrease of $0.0009\pm 0.00177$. In the 100-node experiments, permutating the EXPECTED_COSTGW_OVER_SDP_COST feature lead to a more pronounced decline in performance, measured at $-0.0067\pm 0.0048$. Permutation of the feature LOG_SMALLEST
_EIGVAL results in a smaller drop in performance, recorded as $-0.0004\pm 0.0008$ on performance. Permuting both features together causes a decrease in performance of $-0.0012\pm 0.0013$. Overall, we can conclude that the performance shifts upon permuting a single feature or the top two features slightly affect the model’s performance.
@@
RQAOA as a high-performing heuristic with machine learning pipeline
(a) Partial dependence plot highlighting key features that distinguish significant RQAOA performance gains on 20-node maximum cut instances.
(b) Partial dependence plot highlighting key features that distinguish significant RQAOA performance gains on 100-node maximum cut instances.
[…]
After successfully applying our evolutionary approach to evolve hard maximum cut instances, a new question emerges: whether RQAOA consistently performs well? For our purposes, we seek to establish if RQAOA can achieve a performance level above 0.96, with a minimum advantage of 0.04 over the GW algorithm. We develop classifiers on both 20-node and 100-node instances that incorporate all the features previously discussed with achieved a balanced accuracy of $0.9826\pm 0.0042$ on 20-node experiments and a $0.8343\pm 0.0303$ in the 100-node experiment, maintaining the same experimental and analysis procedures as before. The models are detailed further in section C (see the supplementary material (Pan et al., 2025)). Fig. 6 shows the partial dependent plot for the top two features in both 20-node and 100-node cases, respectively. The calculation of feature importance scores is based on permutation importance (Pedregosa et al., 2011). The analysis of these plots reveals a clear pattern: low values in GW-related features are strongly associated with increased predicted probabilities of RQAOA achieving outstanding performance. Moreover, the significance of the DENSITY feature in influencing RQAOA’s performance is evident. This analysis not only shows the impact of GW-related cost features but also emphasizes the critical role of DENSITY in determining the effectiveness of the RQAOA algorithm.
Table A3 and Table A4 (see the supplementary material (Pan et al., 2025)) showcase the key features identified by the classifier for 20-node and 100-node maximum cut instances, respectively. Highlighted in these tables, the features emphasize the robust performance of RQAOA, validating its effectiveness as a high-performing heuristic through significant permutation importance.
We evaluate the performance of RQAOA against different thresholds to better understand our generated instances. Thresholds of 0.95 and 0.98 were initially considered, but discarded because, in each case, either the 20-node or 100-node instances failed to meet the criteria to be labeled as 1. At a threshold of 0.97, the average balanced accuracy achieved is $0.9795\pm 0.0045$ for 20-node instances, which is comparable to the performance at a threshold of 0.96. LOG_CHROMATIC_NUM and GRAPH_ASSORTATIVITY are identified as the most significant features according to permutation importance calculations. For 100-node instances, the performance improved notably at the 0.97 threshold, achieving an average balanced accuracy of $0.9214\pm 0.0049$. This represents a substantial enhancement over the results obtained with the 0.96 threshold. The most influential features for this larger node size are identified as PERCENT_CLOSE3_LOWER_TRIANGULAR and GRAPH_ASSORTATIVITY. These results demonstrate that the model’s effectiveness varies significantly with different threshold settings, with each threshold providing unique insights, especially in scenarios involving larger graph instances. This responsiveness underscores the need to thoughtfully choose performance benchmarks that align with the unique attributes and requirements of the dataset.
6. Conclusion
Given the current literature does not provide detailed guidelines for constructing hard maximum cut instances, our research addresses this gap. We focus on developing instances that highlight the distinctions between the quantum method RQAOA and the classical GW algorithm by effectively identifying maximum cut instances that pose significant challenges or are notably manageable for the RQAOA compared to the performance of the GW algorithm. Specifically, our work extends beyond situations where RQAOA merely surpasses GW, exploring instances where RQAOA achieves exceptionally high-quality outcomes and significantly outperforms GW. We employ an evolutionary approach using the classical CMA-ES algorithm, enhanced with a fitness function specifically tailored within the latent space of a well-trained Graph Autoencoder.
In predicting these advantages, the machine learning models we developed demonstrate strong results across both 20-node and 100-node instances. It confirms the necessity of both categories of features in analyzing instances where RQAOA significantly outperforms GW, reinforcing the comprehensive nature of the analytical framework. Our findings highlight EXPECTED_COSTGW_OVER_SDP_
COST as the most influential feature when RQAOA is not considered as a high-performing heuristic. This trend remains consistent in 100-node maximum cut instances when we consider RQAOA as a high-performing heuristic but does not hold for 20-node instances. Conversely, DENSITY plays a key role when RQAOA functions as a high-performing heuristic. Beyond these primary features, various other features also contribute to the machine learning classifiers, but GIRTH is found to have no impact in either experimental setting.
Our study offers a fresh perspective on graph generation within network science, encouraging further exploration and advancements in this area. Future work can be explored from various perspectives. Our current experiments focus on graphs with 20 and 100 nodes, but we are actively working to scale up to larger instances, such as 400-node graphs. A key challenge in this expansion is the scalability of the PIGVAE model, originally designed for smaller graphs. Applying it to 400-node graphs causes severe memory bottlenecks. While we’ve reduced the model size to mitigate this, it may impact training performance. Moving forward, we plan to test lightweight GNN architectures and use multi-GPU or distributed training to support larger models. For example, the use of sparse attention mechanisms is worth exploring, such as the implemented in Exphormers (Shirzad et al., 2023). In addition to GNNs, diffusion models present a promising alternative for scalability, which can offer a more efficient scaling path. As a next step, we will focus on 200–400 node graphs to evaluate and refine our approach before scaling further. In parallel, we aim to enrich our analysis by incorporating additional classical heuristics beyond GW, offering a broader perspective on instance difficulty.
Acknowledgments
This work has been supported by the Australian Research Council through grant FT200100536. The authors would like to thank Professor Chunhua Shen from Zhejiang University for insightful discussions on graph neural networks that contributed to the development of this research.
Appendix A Graph Features
The features of the instances utilized in this study are detailed as follows:
  • LOG_NUM_EDGES: The logarithm of the number of edges in the graph.
  • DENSITY: The proportion of actual edges to the maximum possible number of edges (i.e., of a complete graph).
  • LOGRATIO_EDGETONODES: The natural logarithm of the edge-to-node ratio.
  • LOG_CHROMATIC_NUM: The logarithm of the minimum number of colours required to colour the graph such that no two adjacent nodes share the same colour.
  • NORM_MIS: The size of the largest independent node-set, adjusted relative to the total number of nodes, showcasing the proportion of non-adjacent nodes.
  • GRAPH_ASSORTATIVITY: Gauges node connection similarity based on node degree, with values near 1 suggesting similar degree connections.
  • SPECTRAL_GAP: The difference between the largest and second-largest eigenvalues of the graph’s normalized Laplacian matrix.
  • LOG_LARGESTEIGVAL: The logarithm of the highest eigenvalue from the normalized Laplacian matrix.
  • LOG_SECONDLARGESTEIGVAL: The logarithm of the normalized Laplacian matrix’s second-largest eigenvalue.
  • LOG_SMALLESTEIGVAL: The logarithm of the smallest non-trivial eigenvalue from the normalized Laplacian matrix.
  • GIRTH: The length of the shortest loop in the graph.
  • TRANSITIVITY: The likelihood ratio of nodes to form tightly connected clusters.
Appendix B GW-Related Features
The features related to the GW algorithm are introduced as follows:
  • PERCENT_CUT: The ratio of the optimal cut value derived from the semidefinite programming (SDP) solution to the overall number of edges in the graph.
  • PERCENT_POSITIVE_LOWER_TRIANGULAR: Quantifies the share of positive entries in the lower triangular section of the factorized matrix.
  • PERCENT_CLOSE1_LOWER_TRIANGULAR: The proportion of elements in the lower triangular of the factorized matrix that have absolute values under 0.1.
  • PERCENT_CLOSE3_LOWER_TRIANGULAR: The percentage of elements in the lower triangular of the factorized matrix with absolute values smaller than 0.001.
  • EXPECTED_COSTGW_OVER_SDP_COST: The proportion of the GW approach’s average cost (over 1,000 random trials) relative to the maximum theoretical cost as determined by the SDP method.
  • STD_COSTGW_OVER_SDP_COST: Quantifies the standard deviation of the outcomes from 1,000 random implementations of GW, normalized by the best possible cost determined through the SDP method.
Appendix C Machine Learning Classifiers
The models with hyperparameters constructed in the section of feature analysis with the machine learning pipeline are highlighted in the following:
  • The classifier developed using TPOT for instances involving 20 nodes uses an ExtraTreesClassifier, which comprises a collection of 100 decision trees. This ensemble employs the entropy criterion to determine the quality of splits within the trees. Each tree considers approximately 35% of the features for node splitting. It also ensures that each leaf node has at least one sample, while at least nine samples are necessary to initiate further node splitting. The random_state parameter is set to 42, ensuring reproducibility.
  • The classifier built with TPOT for 100-node instances is implemented as an ExtraTreesClassifier that features bootstrapping. The entropy criterion is used to evaluate split quality across the trees. It also requires 80% of the features to be considered for each tree split and a minimum of 19 samples for each leaf node. Additionally, a node split requires at least 5 samples to proceed. The ensemble includes 100 trees, and the random_state parameter is fixed at 42.
The models with hyperparameters constructed in the section of RQAOA as a high-performing heuristic with the machine learning pipeline are highlighted in the following:
  • The classifier configured for 20-node instances using the TPOT library employs a sequential machine-learning pipeline. It begins with a StackingEstimator that deploys an MLPClassifier, characterized by a learning rate of 1.0 and a low regularization factor. This is complemented by another StackingEstimator, which utilizes an ExtraTreesClassifier. This classifier uses the entropy criterion method, avoids bootstrapping, and considers 65% of the features during each split, all while adhering to predefined parameters for sample splits and minimum leaf size. The data is then normalized using a StandardScaler to ensure the features are optimally scaled. Finally, the process culminates with a GaussianNB classifier.
  • The classifier configured using TPOT for 100-node instances also employs a sequential machine-learning pipeline. Initially, a StackingEstimator incorporating a GaussianNB classifier makes preliminary predictions that feed into later stages. Following that, a SelectFwe feature selector uses the f_classif with an alpha of 0.012 to selectively maintain features. Subsequently, another StackingEstimator featuring a Decision Tree Classifier is used, which is tailored with the Gini index for node splitting, a cap of 7 on tree depth, a minimum of 9 samples per leaf, and at least 12 samples required for further splitting. A RobustScaler is then employed to adjust feature scales. The process concludes with a GaussianNB classifier.
Appendix D Important Features with Machine Learning Pipeline
@@
Table A1. Important features with the classifier built on 20-node maximum cut instances.
@@
Table A2. Important features with the classifier built on 100-node maximum cut instances.
Appendix E Important Features with RQAOA as a High-performing Heuristic